### v1

Abstract
Multimodal medical analysis combining image and tabular data has gained increasing attention. However, effective fusion remains challenging due to cross-modal discrepancies in feature dimensions and modality contributions, as well as the noise from high-dimensional tabular inputs. To address these problems, we present AMF-MedIT, an efficient Align-Modulation-Fusion framework for medical image and tabular data integration, particularly under data-scarce conditions. To harmonize dimension discrepancies and dynamically adjust modality contributions, we propose the Adaptive Modulation and Fusion (AMF) module, a novel modulation-based fusion paradigm with a streamlined architecture. We first derive the modulation objectives and introduce a modality confidence ratio, enabling the incorporation of prior knowledge into the fusion process. Then, the feature masks, density and leakage losses are proposed to achieve the modulation objectives. Additionally, we introduce FT-Mamba, a powerful tabular encoder leveraging a selective mechanism to handle noisy medical tabular data efficiently. Furthermore, interpretability studies are conducted to explore how different tabular encoders supervise the imaging modality during contrastive pretraining for the first time. Extensive experiments demonstrate that AMF-MedIT achieves a superior balance between multimodal performance and data efficiency while showing strong adaptability to incomplete tabular data. Interpretability analysis also highlights FT-Mamba’s capabilities in extracting distinct tabular features and guiding the image encoder toward more accurate and flexible attention patterns.
keywords:
[…]
1 Introduction
Rrecently, multimodal learning that integrates medical images and tabular data has emerged as a critical research frontier in precision medicine. While computer-aided diagnosis based on images has been extensively investigated, the complementary role of tabular data in enhancing clinical decision-making is gaining attention. Such tabular data—typically derived from electronic health records (EHRs), laboratory tests, and standardized questionnaires—provide rich clinical context regarding the current symptoms and past medical history, which visual patterns alone cannot capture (Huang et al., 2020). Studies have consistently shown that integrating this clinical context improves image interpretation and supports more accurate diagnoses (Acosta et al., 2022).
However, real-world applications of medical images and tabular data pose two challenges. First, significant disparities in feature dimension and confidence levels across modalities complicate the fusion process. (1) Dimension discrepancy: Image encoders require high-dimensional representations to retain anatomical details, whereas tabular encoders generate more compact features. For example, a ResNet-50 encoder extracts 2,048-dimensional global features, whereas a typical feature vector dimension of TabNet is 128 (Arik & Pfister, 2021). This imbalance in feature dimensions can lead to uneven parameter allocation during fusion, where higher-dimensional features tend to dominate in parameter space. (2) Confidence-level discrepancy: The reliability of each modality in multimodal fusion is influenced by the downstream task and data quality. When images provide definitive pathological signatures but tabular data suffer from missing values, an optimal model is expected to prioritize image evidence. Conversely, in predicting Ischemic Heart Disease (IHD), laboratory and lifestyle factors in tabular data often serve as more reliable prognostic indicators, with images playing a supplementary role (Zambrano Chaves et al., 2023). This disparity in confidence necessitates fusion strategies with adaptive capabilities to appropriately balance multimodal contributions. Second, medical tabular data inherently contain substantial noise, which hinders robust and accurate fusion (Borisov et al., 2024). These data consist of numerous features, many of which can be irrelevant to the target disease. For instance, among the selected 117 data fields for cardiac prediction in (Hager et al., 2023), over 60% have importance scores below 0.005 as measured by integrated gradients. Additionally, the prevalent problem of missing data further amplifies the noise of medical tabular data (Miao et al., 2023).
These challenges, coupled with the persistent data scarcity in the medical domain, place greater demands on both fusion strategies and tabular representation learning, requiring a delicate balance between model capacity and data efficiency. Although the recent adaptation of Self-Supervised Learning (SSL) to the medical field has alleviated data scarcity in encoder training, core challenges in medical image-tabular fusion and tabular feature extraction remain unresolved. Naive concatenation fusion ignores both dimension and confidence disparities, whereas complex cross-attention mechanisms are data-hungry and computationally inefficient. In parallel, while architectures like Mamba (Gu & Dao, 2023) have advanced long-sequence modeling in natural language processing, medical tabular encoding still relies on either simplistic Multi-Layer Perceptrons (MLPs) or transformer-based models that require substantial training data to perform effectively.
In this work, we aim to propose an efficient fusion paradigm that resolves cross-modal disparities through a streamlined architecture, while enhancing the tabular encoder under limited data conditions. Our method is based on three insights. First, inspired by contrastive learning, constructing a unified cross-modal representation space can mitigate representation discrepancies in a straightforward manner. However, unlike matching tasks, medical inference requires preserving modality-specific distinctions (Chen & Hong, 2024). This prompted us to construct a unified yet distinct cross-modal latent space by modulating only the dimension and scale of unimodal features while retaining essential information from the other modality. Second, given the scarcity of medical multimodal data, leveraging prior knowledge from preliminary experiments enables more efficient and interpretable fusion than directly learning cross-modal interactions from limited data. Therefore, we incorporate modality confidence priors into fusion to guide adaptive contribution allocation. Third, medical tabular data can be viewed as long feature sequences containing both relevant and irrelevant elements. Mamba’s selective and sequential modeling capabilities make it well-suited for processing such complex inputs (Gu & Dao, 2023). This motivates us to explore Mamba for medical tabular representation.
Based on the aforementioned insights, we present AMF-MedIT, an efficient Align-Modulation-Fusion framework for integrating medical image and tabular data. In the Align stage, we leverage contrastive learning to train unimodal encoders without annotations. In the Modulation stage, we present a novel paradigm to harmonize cross-modal discrepancies, generating fusion-compatible unimodal representations. Finally, in the Fusion stage, the modulated representations are integrated into a final model for downstream tasks.
Specifically, we propose an Adaptive Modulation and Fusion (AMF) module for unimodal feature modulation. To construct a unified yet distinct latent space for image and tabular features, we first derive the modulation objectives of similar feature densities and adaptive feature dimensions. Meanwhile, we define a modality confidence ratio $R_{conf}$ to incorporate prior information and guide adaptive fusion. In the concrete implementation, we propose feature masks to truncate features and density and leakage losses to compress information, thereby flexibly adjusting the feature dimensions within fixed network structures. Additionally, we introduce FT-Mamba, an adaptation of Mamba that enhances tabular feature extraction. Leveraging the selective State-Space Mechanisms (SSMs), FT-Mamba effectively captures discriminative features from noisy medical records without the need for large-scale training data. Furthermore, we conduct interpretability analysis to observe tabular encoders’ ability to preserve critical indicators and their supervisory role in imaging modality during contrastive learning. Substantial comparative experiments and interpretability studies further demonstrate the effectiveness of our method. The main contributions are summarized as follows.
  • We present AMF-MedIT, an efficient align-modulation-fusion framework for medical image and tabular data. It is designed to address key challenges in the medical domain, particularly in small-sample settings, and features a streamlined modulation-based fusion module coupled with a powerful tabular encoder.
  • We propose the AMF module, a novel modulation-based fusion paradigm that mitigates dimension disparities and adaptively balances modality contributions by incorporating prior knowledge. We introduce FT-Mamba, to the best of our knowledge the first adaptation of Mamba for medical tabular data, aiming to enhance the feature extraction from noisy tabular data.
  • To the best of our knowledge, we conducted the first interpretability study on medical image-tabular contrastive learning, revealing how different tabular encoders impose distinct supervisory patterns on the imaging modality.
  • Experiments on both complete and incomplete tabular data demonstrate the superiority, computational efficiency, and adaptability of our method compared to mainstream fusion methods.
2 Related research
2.1 Self-Supervised Multimodal Learning
Self-supervised learning is proposed to extract meaningful representations from unlabeled data. In multimodal learning, training models still often require expensive human annotation (Krishnan et al., 2022). To mitigate this challenge, SSL has been applied to multimodal settings with various pretext tasks. CLIP utilizes image-text contrastive learning on massive web data, achieving remarkable zero-shot performance (Radford et al., 2021). Subsequent works also combine tasks such as Masked Language Modeling (MLM) and Image-Text Matching (ITM) to enhance cross-modal understanding (Li et al., 2021). Recently, SSL has been explored for image-tabular fusion (Hager et al., 2023; Du et al., 2024; Huang, 2023; Chen & Hong, 2024), but existing studies have not addressed specific challenges in the medical domain. Building on contrastive learning, our approach explicitly tackles the challenges in a medical context.
2.2 Multimodal Medical Image-Tabular Learning
Multimodal image-tabular learning exploits tabular data to facilitate visual task learning, which is widely used in the medical field (Huang et al., 2020). Several studies have employed late fusion, aggregating predictions from separate unimodal models for final decisions (Zambrano Chaves et al., 2023; Zhang et al., 2021). To facilitate deeper and earlier interactions between modalities, joint fusion has become more prevalent. Typically, features from different modalities are concatenated and then processed using fully connected networks (Lee et al., 2023; Vale-Silva & Rohr, 2021), transformer-based (Zheng et al., 2024, 2022), or traditional machine learning models (Zheng et al., 2020) for fusion. To enhance modality interaction, some studies employ cross-attention modules (Jia et al., 2024) to joint unimodal features. However, concatenation-based methods cannot reconcile the discrepancies between unimodal features, resulting in suboptimal fusion performance. Meanwhile, cross-attention methods rely entirely on full pairwise attention without leveraging prior knowledge, leading to high computational overhead, and making it difficult to achieve effective learning on small-scale medical datasets.
While SSL has advanced image-tabular learning, fusion strategies remain unchanged. Most studies still rely on feature concatenation (Hager et al., 2023; Huang, 2023; Painchaud et al., 2024) or cross-attention mechanisms (Zhou et al., 2023; Du et al., 2024; Chen & Hong, 2024; Chen et al., 2024). To address this limitation, we propose a novel fusion paradigm that flexibly regulates unimodal features within a streamlined structure, optimizing the trade-off between fusion performance and data efficiency.
2.3 Deep Learning for Tabular Data
Tabular data present unique challenges for deep learning (DL) models, requiring them to handle various data-related issues such as noise, imprecision, missing values, diverse attribute types, and varying value ranges (Borisov et al., 2024). Existing research on deep learning for tabular data provides diverse novel architectures. Gorishniy et al. (Gorishniy et al., 2021) compared the main deep architectures, including MLP, ResNet, and FT-Transformer, finding ResNet to be a simple yet effective baseline, whereas FT-Transformer serves as a universal architecture. Recently, Mamba has gained attention for leveraging selective SSMs to achieve linear-time complexity and memory efficiency (Gu & Dao, 2023). Thielmann et al. explored its application in the tabular domain, demonstrating Mamba’s feasibility for tabular data (Thielmann et al., 2024).
Despite this progress, tabular data processing in the medical domain remains relatively underdeveloped. Some studies use raw tabular data as input (Zheng et al., 2020; Pölsterl et al., 2021), while others employ MLP-like architectures for simple processing (Lee et al., 2023; Zheng et al., 2022; Vale-Silva & Rohr, 2021; Hager et al., 2023). However, such simplistic models often fall short of capturing meaningful representations from noisy and high-dimensional medical tabular data. Recent works have adopted transformer-based models to enhance tabular feature extraction (Zheng et al., 2024; Zhou et al., 2023; Du et al., 2024; Painchaud et al., 2024; Chen & Hong, 2024; Chen et al., 2024). Nevertheless, these models require large datasets for effective training and often encounter optimization difficulties on limited medical data. Inspired by FT-Transformer and Mambular, we propose a streamlined adaptation of the Mamba architecture for tabular data, specifically designed to handle the noisy nature of medical tabular data.
3 Methods
3.1 Overall Framework
Figure 1: The overall framework of AMF-MedIT. Modules outlined with black dashed borders are being trained in each stage. We employ FT-Mamba, a powerful and efficient encoder, for medical tabular data. Additionally, the AMF module is designed to modulate unimodal features, thereby facilitating effective fusion. Notably, $R_{conf}$ can be flexibly adjusted based on downstream requirements.
The overall framework of AMF-MedIT is shown in Fig. 1. The training pipeline comprises two stages: (a) Imaging-tabular pretraining and (b) Fine-tuning for adaptive modulation and fusion. In each stage, the image-tabular pairs, ($\boldsymbol{x}_{img}\in\mathbb{R}^{H\times W\times D}$, $\boldsymbol{x}_{tab}\in\mathbb{R}^{N}$) are fed into the corresponding unimodal encoders, $\boldsymbol{E}_{img}$ and $\boldsymbol{E}_{tab}$. Here, $H$, $W$, and $D$ denote the height, width, and number of channels of the input image, respectively, and $N$ represents the number of tabular features. We use ResNet-50 (He et al., 2016) as the image encoder and FT-Mamba as the tabular encoder, which will be introduced in Section 3.2. The encoders then generate the original unimodal embeddings, $\boldsymbol{z}_{img}\in\mathbb{R}^{D_{img}}$ and $\boldsymbol{z}_{tab}\in\mathbb{R}^{D_{tab}}$, where $D_{img}$ and $D_{tab}$ denote the dimensions of unimodal features.
In the pretraining stage, the projection heads $\boldsymbol{\Phi}_{img}$ and $\boldsymbol{\Phi}_{tab}$ map $\boldsymbol{z}_{img}$ and $\boldsymbol{z}_{tab}$ into a shared latent space as $\boldsymbol{\hat{z}}_{img}\in\mathbb{R}^{D_{\Phi}}$ and $\boldsymbol{\hat{z}}_{tab}\in\mathbb{R}^{D_{\Phi}}$ separately. $D_{\Phi}$ denotes the dimensions of projections. On the supervise of the Image-Tabular Contrastive (ITC) loss, the projections are pulled and pushed to maximize the similarity between matching image-tabular pairs while minimizing the similarity of non-matching pairs.
In the fine-tuning stage, we propose a simple AMF module for multimodal fusion. Guided by the proposed density and leakage losses, this module transforms the original unimodal features, $\boldsymbol{z}_{img}$ and $\boldsymbol{z}_{tab}$, into modulated features, $\boldsymbol{\tilde{z}}_{img}$ and $\boldsymbol{\tilde{z}}_{tab}$, which are then combined to form a cohesive fused feature, $\boldsymbol{z}_{fused}$. Finally, $\boldsymbol{z}_{fused}$ is fed to a Fully-Connected (FC) layer to generate the classification probability.
3.2 FT-Mamba
Following the design of FT-Transformer (Gorishniy et al., 2021), we adapt the Mamba architecture (Gu & Dao, 2023) for tabular feature extraction and name it as FT-Mamba (Feature Tokenizer + Mamba). As shown in Fig. 1 (a), we first separately tokenize the categorical and numerical tabular features to embeddings and add a [CLS] token at the sequence end. Then a stack of mamba blocks is used to extract the high-dimensional features from the embeddings, and the final representation of the [CLS] token is used as the original tabular embedding, $\boldsymbol{z}_{tab}$.
Figure 2: The key modules in FT-Mamba. (a) Feature tokenizer: categorical and numerical tabular features are separately tokenized into embeddings; (b) a Mamba block, and its core module (c) Selective SSM: it enables input-dependent parameterization of SSMs and dynamically adjusts them to focus on relevant input features.
3.2.1 Feature Tokenizer
Our feature tokenizer module follows the design of (Gorishniy et al., 2021). It transfers the tabular input $\boldsymbol{x}_{tab}$ to embeddings $\boldsymbol{t}_{tab}\in\mathbb{R}^{N\times D_{tab}}$ as shown in Fig. 2 (a). For continuous features, the feature value is mapped by a linear function with $D_{tab}$-dimensional weights and bias. For categorical features, each category is mapped to a vector through a lookup table and is added with the corresponding bias.
3.2.2 Mamba Blocks
As shown in Fig. 2 (b) and (c), Mamba maps input token ${\boldsymbol{t}_{tab}}$ to the representation ${\boldsymbol{z}_{tab}}$ through a hidden state ${\boldsymbol{h}_{tab}}$, which stores and updates long-term information without explicit self-attention. Moreover, the module employs adaptive parameters $\mathbf{B},\mathbf{C}$ as functions of the input, allowing it to selectively focus on relevant information while filtering out irrelevant features. Unlike transformers, which compute all pairwise interactions, Mamba reduces unnecessary computations, facilitating efficient optimization in data-limited scenarios.
3.3 Adaptive Modulation and Fusion Module
3.3.1 Problem Formulation
To address the cross-modal discrepancies in feature dimensionality and confidence levels, we first quantify the characteristics of unimodal features. Since the features have been processed by the deep encoders, redundant or homogeneous information with low entropy has already been discarded. Therefore, we use the $\ell_{1}$-norm, which intuitively reflects the sparsity of feature vectors, to measure the information content of a unimodal feature as in (1). It reflects the feature’s informational contribution to the classification decision.
[…]
Then we the define feature density of a feature vector as the ratio of its content to feature dimension, as shown in (2). It describes the average information content on each element and reflects the overall amplitude. Equation (1) and (2) use $\boldsymbol{z}_{img}$ as an example, and the same concept applies to the tabular modality.
[…]
Additionally, we introduce the modality confidence ratio $R_{\text{conf}}$ to guide the reliability of each modality in downstream tasks. This ratio is defined as the confidence level of image features relative to tabular features in the current downstream task, as shown in (3). In our experiments, $R_{\text{conf}}$ is calculated by the unimodal performance of pretrained encoders.
[…]
Next, we propose the quantitative requirements for the unified yet distinct latent space. To prevent features with large-scale values from overwhelming those with small-scale values, the feature densities across modalities should be similar, as formulated in (4a). Given comparable densities, the information content of each modality can be directly reflected by the feature dimensions according to (2), enabling flexible adjustment of modality contributions by modulating corresponding feature dimensions. Ideally, modality contributions in fusion should align with their confidence levels. Let $L_{img}$ and $L_{tab}$ denote the dimension of $\boldsymbol{\tilde{z}}_{img}$ and $\boldsymbol{\tilde{z}}_{tab}$, respectively, this requirement is formulated in (4b).
$\textbf{Density}(\boldsymbol{\tilde{z}}_{img})$ $\approx\textbf{Density}(\boldsymbol{\tilde{z}}_{tab})$ (4a)
$\frac{\textbf{Confidence}(\boldsymbol{\tilde{z}}_{img})}{\textbf{Confidence}(\boldsymbol{\tilde{z}}_{tab})}$ $\approx\frac{L_{img}}{L_{tab}}\approx R_{\text{conf}}$ (4b)
According to the objective outlined in (4), we propose the AMF module to modulate the original unimodal features, which constrains similar feature densities across modalities while adjusting the feature dimensions based on their confidence levels. It employs a straightforward architecture and simple training tasks to achieve accurate and efficient fusion.
3.3.2 Implementation
The structure of the AMF module is shown in the Fig. 3. First, two projector heads, $\boldsymbol{\Psi}_{img}$ and $\boldsymbol{\Psi}_{tab}$, which consistent with those used in contrastive pretraining, map the original features $\boldsymbol{z}_{img}$ and $\boldsymbol{z}_{tab}$ to the transitional features $\boldsymbol{z}\textquoteright_{img}\in\mathbb{R}^{D_{out}}$ and $\boldsymbol{z}\textquoteright_{tab}\in\mathbb{R}^{D_{out}}$. Here, $D_{out}$ denotes the input dimension of the final model. Then, a set of feature masks, leakage loss, and density loss are proposed to achieve the objective defined in (4), resulting in the flexible-dimension modulated features $\boldsymbol{\tilde{z}}_{img}$ and $\boldsymbol{\tilde{z}}_{tab}$. Finally, $\boldsymbol{\tilde{z}}_{img}$ and $\boldsymbol{\tilde{z}}_{tab}$ are concatenated to obtain the fused feature $\boldsymbol{z}_{fused}$.
Figure 3: The structure of the AMF module. For each modality, the original features are mapped to the transitional features separately. Then feature masks implicitly partition them into retained parts (solid lines) and discarded parts (dashed lines). During training, the leakage loss compresses effective information into the retained parts, and density loss aligns their density across modalities.
Feature mask.
One challenge in our framework is how to flexibly adjust the dimension of the modulated unimodal features to fit the fixed architecture of the final model, which receives the concatenated feature as input and expects a specific dimension $D_{out}$. That is, the dimensions of the modulated features, $L_{img}$ and $L_{tab}$, are constrained by $D_{out}$ as follows:
$L_{img}+L_{tab}=D_{out}.$ (5)
However, $L_{img}$ and $L_{tab}$ should vary with the $R_{conf}$ to regulate the modality contribution according to (4b). By combining (4b) and (5), specific values of $L_{img}$ and $L_{tab}$ can be derived as follows:
$\begin{cases}L_{img}&=D_{out}-\left\lfloor\frac{D_{out}}{1+R_{conf}}\right\rfloor\\ L_{tab}&=\left\lfloor\frac{D_{out}}{1+R_{conf}}\right\rfloor\end{cases}.$ (6)
Therefore, we defined a feature mask for each modality. For the imaging modality, the mask is a binary vector of dimension $D_{out}$, with the first $L_{img}$ elements set to $1$ and the rest set to 00; Conversely, the tabular feature mask is defined in the opposite way as follows:
$\begin{cases}\boldsymbol{m}_{img}&=[\underbrace{1,1,\dots,1}_{L_{img}},\underbrace{0,0,\dots,0}_{L_{tab}}]^{T}\\ \boldsymbol{m}_{tab}&=[\underbrace{0,0,\dots,0}_{L_{img}},\underbrace{1,1,\dots,1}_{L_{tab}}]^{T}\end{cases}.$ (7)
These masks divide the transitional features into a retained part and a discarded part. The retained parts are concatenated to form the fused feature, while the discarded parts which contain little information are dropped as follows:
$\boldsymbol{z}_{fused}={\boldsymbol{m}_{img}}^{T}\boldsymbol{z}\textquoteright_{img}+{\boldsymbol{m}_{tab}}^{T}\boldsymbol{z}\textquoteright_{tab}.$ (8)
Leakage loss.
To prevent the discarded part from carrying away crucial information, we introduce a leakage loss that quantifies the information density in the masked-out region, as defined in (9), where $\mathcal{N}$ is the batch size.
$\mathcal{L}_{\text{leakage}}=\frac{1}{2\mathcal{N}}\sum_{j\in\mathcal{N}}\left(\frac{{\left\|{\boldsymbol{m}_{tab}}^{T}{\boldsymbol{z}\textquoteright_{img}}^{j}\right\|}_{1}}{D_{tab}}+\frac{{\left\|{\boldsymbol{m}_{img}}^{T}{\boldsymbol{z}\textquoteright_{tab}}^{j}\right\|}_{1}}{D_{img}}\right)$ (9)
By minimizing the leakage and the cross-entropy (CE) losses, both $\boldsymbol{\Psi}_{img}$ and $\boldsymbol{\Psi}_{tab}$ are trained to compress effective information into the retained part, preventing the adjustment of feature dimension from reducing feature effectiveness.
Density loss.
To align the density of the retained features as per (4a), we introduce a density loss, defined as the density difference between unimodal features in (10). By minimizing the density loss, we encourage the modulated features to exhibit similar density.
$\mathcal{L}_{\text{density}}=\frac{1}{\mathcal{N}}\sum_{j\in\mathcal{N}}\left|\frac{{\left\|{\boldsymbol{m}_{img}}^{T}{\boldsymbol{z}\textquoteright_{img}}^{j}\right\|}_{1}}{D_{img}}-\frac{{\left\|{\boldsymbol{m}_{tab}}^{T}{\boldsymbol{z}\textquoteright_{tab}}^{j}\right\|}_{1}}{D_{tab}}\right|$ (10)
3.4 Optimization
In the pretraining stage, image-tabular pairs are used for contrastive learning and the ITC loss is applied (Hager et al., 2023). The image-to-tabular contrastive loss is defined in (11) and the tabular-to-image contrastive loss shares a symmetric formulation.
$\mathcal{L}_{\text{i2t}}=-\frac{1}{\mathcal{N}}\sum_{j\in\mathcal{N}}log\frac{\text{exp}(\text{cos}({\boldsymbol{\hat{z}}_{img}}^{j},{\boldsymbol{\hat{z}}_{tab}}^{j})/\tau)}{\sum\limits_{k\in\mathcal{N},k\neq j}\text{exp}(\text{cos}({\boldsymbol{\hat{z}}_{img}}^{j},{\boldsymbol{\hat{z}}_{tab}}^{k})/\tau)}\$ (11)
Here, $\mathcal{N}$ represents the batch size, and $\tau$ is the temperature hyperparameter. The overall ITC loss is computed as the average of these two components as: $\mathcal{L}_{\text{ITC}}=\left(\mathcal{L}_{\text{i2t}}+\mathcal{L}_{\text{t2i}}\right)/2.$
In the fine-tuning stage, the loss function consists of three components. Alongside the leakage loss in (9) and density loss in (10) for multimodal fusion, we employ CE loss for classification. For the OL3I dataset, which suffers from severe class imbalance, we apply a weighted CE loss as defined in (12), where $p$ is the model output, $y$ is the classification label, and $\alpha$ is a hyperparameter for mitigate class imbalance.
$\mathcal{L}_{\text{weighted ce}}=-\frac{1}{\mathcal{N}}\sum_{j\in\mathcal{N}}\left(\alpha y^{j}log(p^{j})-(1-\alpha)(1-y^{j})log(1-p^{j})\right)$ (12)
For the DVM dataset, standard CE loss is used as formulated in (13), where $\mathcal{C}$ denotes the number of classes.
$\mathcal{L}_{\text{ce}}=-\frac{1}{\mathcal{N}}\sum_{j\in\mathcal{N}}\sum_{c\in\mathcal{C}}y_{c}^{j}log(p_{c}^{j})$ (13)
The overall loss function is a weighted sum of these three losses, with $\lambda_{1},\lambda_{2}$ as hyperparameters to balance leakage and density loss as follows:
$\mathcal{L}_{\text{OL3I}}$ $=\mathcal{L}_{\text{weighted ce}}+\lambda_{1}\mathcal{L}_{\text{leakage}}+\lambda_{2}\mathcal{L}_{\text{density}}$ (14a)
$\mathcal{L}_{\text{DVM}}$ $=\mathcal{L}_{\text{ce}}+\lambda_{1}\mathcal{L}_{\text{leakage}}+\lambda_{2}\mathcal{L}_{\text{density}}.$ (14b)
4 Experiments
4.1 Datasets
We conducted experiments on two multimodal datasets: OL3I (Opportunistic L3 computed tomography slices for Ischemic Heart Disease risk assessment) (AIMI, 2023) and DVM (Data Visual Marketing) (Huang et al., 2021). OL3I is a medical dataset that combines grayscale computed tomography (CT) images with high-dimensional clinical tabular features, while DVM is a large-scale natural dataset for automotive applications, offering a complementary benchmark for general-purpose scenarios.
The OL3I dataset contains 8,139 axial CT slices at the third lumbar vertebra (L3) level, each paired with structured medical record data. Every image is a 512$\times$512 grayscale slice, and the tabular modality includes 423 features covering patient demographics, laboratory results, and aggregated clinical codes. In our study, we adopted the 1-year IHD prediction task as the downstream objective. Due to the severe class imbalance, we used the Area Under the receiver operating Characteristic Curve (AUC) as the evaluation metric. All experiments followed the dataset’s predefined training, validation, and test splits.
The DVM dataset comprises 35,719 image-tabular pairs, each consisting of a 300$\times$300 front-view RGB image of a car and corresponding tabular data on its sales and technical data. To avoid label leakage in the car model classification task, we excluded features such as brand, model year, and vehicle size, resulting in 13 tabular input features. Following the protocol in (Hager et al., 2023), we considered 101 target classes, each with at least 100 samples. Top-1 accuracy (ACC) was used as the evaluation metric, and the data was randomly split into training, validation, and test sets with a ratio of 72%, 18%, and 10%, respectively. Additional preprocessing details are provided in Appendix A.1.
4.2 Experimental Settings
4.2.1 Evaluating Tabular Encoders in Supervised Learning
We first compared four tabular encoders, FT-Mamba, MLP (Hager et al., 2023), ResNet (Gorishniy et al., 2021), and FT-Transformer (Gorishniy et al., 2021), by training unimodal tabular models under supervised learning. Each model consisted of the corresponding encoder followed by an FC classification layer. The number of encoder blocks was fixed to 2 across all models, following the configurations in (Hager et al., 2023; Gorishniy et al., 2021). Based on hyperparameter search, the tabular embedding dimension $D_{tab}$ was set to 256 for MLP and ResNet, and 64 for FT-Transformer and FT-Mamba. Results are reported as the mean and standard deviation across five different random seeds. Additional implementation details are provided in Appendix A.3.
To investigate how the tabular encoders process high-dimensional tabular inputs, we performed Integrated Gradient (IG) analysis on the OL3I dataset. IG estimates the contribution of each input feature by integrating gradient values along a straight-line path from a baseline input to the target sample (Sundararajan et al., 2017). In our setting, the baseline input was defined as the mean of all training tabular samples. IG scores were computed for all training samples with respect to the disease class, and the final feature importance values were averaged to mitigate randomness. For consistent comparison, each feature’s importance was normalized as a percentage of the total importance within the model, thereby highlighting its relative influence during inference.
4.2.2 Multimodal Contrastive Pretraining and Interpretation
We selected ResNet-50 as the image encoder and pretrained it in combination with each of the four tabular encoders introduced in Section 4.2.1, resulting in four distinct image-tabular encoder pairs. All models were trained from scratch using contrastive learning. An exception was made for the OL3I experiment involving ResNet-50 and FT-Transformer, where the contrastive loss failed to converge; in this case, we initialized the ResNet-50 with weights from the ResNet-50-MLP pretraining. Additional implementation details are provided in Appendix A.2.
To investigate how different tabular encoders influence the image modality under contrastive learning, we applied Grad-CAM (Selvaraju et al., 2017) to generate class activation maps (CAMs) from the pretrained multimodal models. For a consistent comparison, all models used a concatenation-based fusion framework to avoid confounding effects from fusion modules.
4.2.3 Multimodal Fine-tuning for Downstream Tasks
In the fine-tuning stage, we compared three multimodal fusion methods: concatenation, cross-attention, and the proposed AMF module. In the concatenation fusion method, we followed the approach in (Hager et al., 2023), where the unimodal feature vectors generated by the pretrained encoders were concatenated and fed into an FC layer for classification. In the cross-attention fusion method, we followed the approach in (Du et al., 2024). The feature map from the final layer of ResNet-50 was flattened into image tokens. These tokens, along with the tabular token generated by FT-Transformer or FT-Mamba, were then fed into a multimodal cross-attention module with 2 layers and an embedding dimension of 512. Finally, the CLS token was passed to the classification layer for prediction. In the AMF fusion, the configurations of modality confidence ratio $R_{conf}$ are provided in Appendix A.4.
We applied both frozen and trainable strategies to each pretrained model during fine-tuning. In the frozen strategy, all parameters of the pretrained encoders were frozen. In the retraining strategy, all weights in the model were trainable. A unimodal ResNet-50 was also trained with supervised learning for comparison. More parameter settings can be found in Appendix A.3.
4.2.4 Experiments on Incomplete Tabular Data
Incomplete tabular data is a common challenge in real-world clinical applications. This variability in downstream settings necessitates that robust multimodal models remain effective even when one modality is partially degraded. To evaluate model adaptability under such conditions, we simulated an incomplete tabular input scenario on the OL3I dataset by masking the most important tabular feature, which was replaced with its mean value across all samples. Feature importance was determined using IG analysis based on the supervised tabular models.
Under this incomplete setting, we fine-tuned the multimodal models using a ResNet-50 image encoder and an FT-Mamba tabular encoder, both pretrained via contrastive learning. Three fusion strategies—feature concatenation, cross-attention, and the proposed AMF module with different $R_{conf}$ values—were compared to assess their robustness to missing tabular information.
4.2.5 Ablation Study on the AMF Module
To validate the necessity of dynamically adjusting unimodal feature dimensions based on modality confidence, we conducted an ablation study on the modality confidence ratio $R_{conf}$ in the AMF module. We tested $R_{conf}$ values of 1.0, 0.97, 0.95, and 0.1 during the fine-tuning on the OL3I dataset, which directly controls the allocation of feature dimensions between the image and tabular modality prior to concatenation. Among these, 1.0 and 0.97 correspond to the $R_{conf}$ values selected for ResNet and MLP tabular encoders, respectively, while 0.1 was chosen to simulate a scenario where the dimensionality of tabular features is significantly higher than that of the image features. This setting contrasts with the concatenation-based fusion, where $R_{conf}$ is not applied.
To evaluate the contribution of the proposed auxiliary loss functions—density loss and leakage loss, we fixed $R_{conf}=0.95$ and compared four different loss configurations during fine-tuning: (1) $\mathcal{L}_{\text{fine-tuning}}=\mathcal{L}_{\text{ce}}$; (2) $\mathcal{L}_{\text{fine-tuning}}=\mathcal{L}_{\text{ce}}+\lambda_{1}\mathcal{L}_{\text{leakage}}$; (3) $\mathcal{L}_{\text{fine-tuning}}=\mathcal{L}_{\text{ce}}+\lambda_{2}\mathcal{L}_{\text{density}}$; (4) $\mathcal{L}_{\text{fine-tuning}}=\mathcal{L}_{\text{ce}}+\lambda_{1}\mathcal{L}_{\text{leakage}}+\lambda_{2}\mathcal{L}_{\text{density}}$ (ours).
4.3 Unimodal Performance of Tabular Models
4.3.1 Performance
Table 1: Performance of unimodal tabular models trained with supervised learning.
Encoder OL3I AUC (%) DVM ACC (%)
MLP 74.67 $\pm$ 2.06 85.50 $\pm$ 0.50
ResNet 74.43 $\pm$ 1.08 85.19 $\pm$ 0.37
FT-Transformer 74.66 $\pm$ 1.25 94.30 $\pm$ 0.29
FT-Mamba 77.36 $\pm$ 1.58 92.71 $\pm$ 0.28
The unimodal performance of the four tabular encoders under supervised training is shown in Table 1, with the best results highlighted in bold. Overall, FT-Mamba demonstrates strong performance on both datasets, particularly excelling on the medical dataset, while FT-Transformer shows a clear advantage only on the large-scale DVM dataset. On the medical OL3I dataset, FT-Mamba significantly outperforms the other encoders, achieving an average AUC of 77.36%. On the DVM dataset, which contains a larger number of samples, FT-Transformer achieves the highest average accuracy of 94.30%, followed closely by FT-Mamba with 92.71%. In contrast, MLP and ResNet perform substantially worse on both datasets.
4.3.2 Integrated Gradient Analysis
(a) MLP
(b) ResNet
(c) FT-Transformer
(d) FT-Mamba
Figure 4: Feature importance distribution for unimodal tabular models on the OL3I dataset.
(a) Importance proportions by category
(b) High-importance features of FT-Mamba
(c) High-importance features of FT-Transformer
Figure 5: IG analysis results for unimodal tabular models. (a): Feature importance proportions by category for the OL3I dataset. Features with an importance percentage less than 1% were classified as low-importance features (light blue). Features with an importance percentage of 1% or greater were classified as high-importance features. Based on (Zambrano Chaves et al., 2023) and established medical knowledge, we further divided the high-importance features into medically relevant features (pink) and low-relevance features (dark blue). (b) and (c): High-importance features and their importance proportion for FT-Mamba and FT-Transformer. (exp_wt_lab_systolic: exponentially weighted average of systolic blood pressure. CPT: Current Procedural Terminology codes, ATC: Anatomical Therapeutic Chemical codes.)
Our IG analysis on the OL3I dataset further interprets the behavior of the four tabular encoders. Fig. 4 shows the overall distribution of feature importance across all tabular features. Due to the high dimensionality of the tabular data, we additionally grouped the 423 features into three categories and visualized their respective importance distributions in Fig. 5(a). More detailed IG results are available in Appendix B.
The results indicate that MLP and ResNet exhibit limited capacity for distinguishing tabular features. As illustrated in Fig. 4(a), MLP tends to distribute its attention relatively uniformly across all input features. Although ResNet demonstrates a single peak in its importance distribution, Table B.1 shows that its most important features account for only 18% of the total importance—substantially lower than the 45% observed in FT-Mamba and the 36% in FT-Transformer. Furthermore, as shown in Fig. 5(a), more than 60% of ResNet and MLP’s attention is dispersed across a large number of low-importance features.
In contrast, FT-Mamba and FT-Transformer are more effective at identifying meaningful features from high-dimensional tabular data. Over 60% of their attention is highly concentrated on a small subset of high-importance features, while other features receive minimal attention. This strong feature discrimination capacity contributes to their superior performance when processing high-dimensional inputs.
However, further comparing the FT-Mamba and FT-Transformer, the high-importance but low-relevance-in-medicine features of FT-Transformer account for 10.63%, which is higher than FT-Mamba and even ResNet. The specific high-importance features for them are presented in Fig. 5(b) and 5(c). All of the high-importance features for FT-Mamba are strongly medically relevant to IHD, including cardiovascular risk factors such as age, CCI, and sex, which have been validated as the most important features for IHD prediction in (Zambrano Chaves et al., 2023), along with other records related to the cardiovascular system. In contrast, FT-Transformer places significant attention on several irrelevant features, such as A12, A06, and J05. This erroneous focus shows that FT-Transformer captures several noise patterns when the training data is limited, which may be the factor for the inferior performance of FT-Transformer compared to FT-Mamba on the OL3I dataset.
4.4 Multimodal Model Evaluation on Complete Tabular Data
4.4.1 Model Interpretation in Contrastive Learning
Fig. 6 shows the CAMs generated by four pretrained image-tabular encoder pairs, along with a unimodal ResNet-50 trained using supervised learning. The CAMs generated by the supervised ResNet-50 lack distinct focal regions, indicating insufficient supervision and limited ability of the image encoder to capture meaningful visual features. CAMs from models using MLP or ResNet as tabular encoders show concentrated attention around the lumbar spine, exhibiting a localized and singular focused pattern. In contrast, models with FT-Mamba and FT-Transformer encoders show a more flexible, multi-point attention distribution.
Figure 6: CAMs generated by supervised ResNet-50 and pretrained multimodal models for IHD prediction. Group (a) illustrates activation patterns in non-IHD cases, while group (b) are IHD-positive samples.
These findings are consistent with the tabular IG analysis in Section 4.3.2 and reflect how different tabular encoders guide the image modality in contrastive learning. ResNet-50, as a convolutional network, inherently emphasizes local feature extraction and is limited in global modeling. When paired with relatively weak tabular encoders like MLP and ResNet, the supervision signal is limited, resulting in minimal changes in the spatial focus of the image encoder. In contrast, FT-Transformer and FT-Mamba provide stronger global supervision through their advanced long-range modeling capacities, enabling the image encoder to learn more structured and informative visual representations.
However, the model with FT-Transformer exhibits suboptimal attention, with substantial activation over irrelevant or blank regions, suggesting that its noisier feature patterns misdirected the image encoder during contrastive training. In contrast, the model with FT-Mamba effectively attends to clinically relevant anatomical structures, including the surrounding musculature, subcutaneous and visceral fat, and cardiovascular regions. Notably, visceral fat proportion is a key biomarker in predicting IHD (Zambrano Chaves et al., 2023). These results highlight FT-Mamba’s superior capability to support informative feature learning in the image modality, especially under the limited data conditions of the OL3I dataset.
4.4.2 Multimodal Performance on Downstream Classification
The performance of the multimodal models on the test set is summarized in Table 2. All models utilizing trainable fusion strategies substantially outperform the unimodal ResNet-50 trained solely on image data. Among all combinations of fusion methods and tabular encoders, our proposed model—AMF module combined with FT-Mamba—achieves the highest AUC scores on the OL3I dataset under both frozen and trainable strategies (76.97% and 78.36%, respectively). On the large-scale DVM dataset, the best results are obtained by the cross-attention + FT-Mamba model in the frozen setting (96.96% ACC), and by the AMF + FT-Transformer model in the trainable setting (99.23% ACC).
Table 2: Classification performance of multimodal methods on complete tabular data.
Method OL3I AUC (%) DVM ACC (%)
[…]
Concatenation MLP 76.31 $\pm$ 0.45 77.18 $\pm$ 0.82 93.51 $\pm$ 1.06 97.45 $\pm$ 0.19
ResNet 76.38 $\pm$ 0.60 74.15 $\pm$ 1.76 92.20 $\pm$ 0.92 97.25 $\pm$ 0.53
FT-Transformer 75.51 $\pm$ 0.27 75.74 $\pm$ 0.81 85.40 $\pm$ 2.33 96.10 $\pm$ 0.39
FT-Mamba 75.80 $\pm$ 0.24 76.49 $\pm$ 0.60 89.08 $\pm$ 1.06 97.41 $\pm$ 0.36
Cross-attention FT-Transformer 75.71 $\pm$ 0.39 75.64 $\pm$ 0.34 93.53 $\pm$ 0.45 97.30 $\pm$ 0.77
FT-Mamba 73.36 $\pm$ 0.30 73.48 $\pm$ 0.98 96.96 $\pm$ 0.52 98.15 $\pm$ 0.31
AMF Module MLP 76.63 $\pm$ 0.22 76.86 $\pm$ 0.35 94.92 $\pm$ 0.43 97.88 $\pm$ 0.42
ResNet 76.50 $\pm$ 0.83 75.46 $\pm$ 0.92 95.09 $\pm$ 0.54 97.28 $\pm$ 0.70
FT-Transformer 75.62 $\pm$ 0.49 76.46 $\pm$ 1.57 93.83 $\pm$ 0.90 99.23 $\pm$ 0.35
FT-Mamba 76.97 $\pm$ 0.75 78.36 $\pm$ 0.12 94.76 $\pm$ 0.45 98.89 $\pm$ 0.11
Supervised ResNet-50 58.29 $\pm$ 4.16 93.00 $\pm$ 1.05
  • 1
    Frozen strategy, where the parameters of pretrained encoders did not participate in gradient updates.
  • 2
    Trainable strategy, where all weights in the model were trainable.
In the concatenation-based fusion, models using FT-Transformer and FT-Mamba underperform those using MLP and ResNet on both datasets. This contrasts with the superior performance of sequence-based encoders in unimodal training and contrastive pretraining. The performance drop can be attributed to the severe dimensional discrepancy between tabular features generated by sequence models ($D_{tab}=64$) and image features ($D_{img}=2048$). Sequence-based tabular encoders produce more compressed and lower-dimensional representations than MLP and ResNet, placing the tabular modality at a particular disadvantage during naive concatenation, where no interaction mechanism exists to compensate for disparities in feature scale and informativeness.
In contrast, the AMF module mitigates this imbalance by modulating unimodal features to comparable dimensions, restoring the advantages of sequence models in multimodal fusion. Compared to concatenation, the AMF module significantly improves performance, with an AUC increase of up to 1.87 percentage points on the OL3I dataset (with FT-Mamba) and an accuracy improvement of up to 8.43 percentage points on the DVM dataset (with FT-Transformer). When paired with MLP or ResNet tabular encoders, the AMF module also consistently outperforms concatenation across most settings.
Comparing the AMF module with the cross-attention module, the AMF module demonstrates a clear advantage under limited data conditions on the OL3I dataset. With FT-Mamba as the tabular encoder, AMF improves the AUC by 3.61 and 4.88 percentage points under frozen and trainable strategies, respectively. On the DVM dataset, where data is abundant, AMF performs comparably to the cross-attention module, suggesting that both methods are capable of effective representation learning when sufficient training samples are available.
Further comparison across both datasets reveals that AMF’s advantage over cross-attention is more prominent under the trainable strategy. This indicates that AMF is more responsive to encoder updates and better leverages the evolving feature representations during training. In contrast, the cross-attention module, with its complex and conservative full-attention structure, tends to be more stable in scenarios with frozen encoders and ample data. However, under the trainable strategy, its increased parameter complexity imposes optimization challenges, which hinders its effectiveness even on the DVM dataset.
4.4.3 Computational Efficiency of Fusion Modules
To compare the computational efficiency of the cross-attention and the proposed AMF module, we analyzed their number of parameters, floating-point operations (FLOPs), and inference latency in Table 3. All measurements were conducted with a batch size of 1. The inference latency was measured on an NVIDIA RTX 4090 GPU, averaged over 100 runs.
Table 3: Comparison of parameter count, FLOPs, and inference latency of the cross-attention and AMF modules.
Fusion Module #Param (M) OL3I DVM
FLOPs(GMac) Latency (ms) FLOPs(GMac) Latency (ms)
Cross-attention 9.49 3.67 1.54 3.17 1.21
AMF Module 8.53 0.0085 0.12 0.0085 0.12
  • 1 GMac = $10^{9}$ multiply–accumulate operations.
  • The input image resolutions are 512 $\times$ 512 for the OL3I dataset and 128 $\times$ 128 for the DVM dataset.
Notably, the cross-attention module exhibits significantly higher parameter counts, FLOPs, and slower inference speed compared to the proposed AMF module, indicating greater model capacity and computational complexity. While such high-capacity architectures may enhance performance on large-scale datasets, they are also more susceptible to overfitting in low-data regimes, as they tend to memorize spurious correlations. In contrast, the lightweight AMF module offers a more streamlined architecture that generalizes better under data-scarce scenarios.
Importantly, the computational cost of cross-attention increases quadratically with the spatial resolution of input images, since its attention mechanism operates over the entire image patch sequence. This sensitivity to image size can greatly amplify both FLOPs and inference time under high-resolution settings. In contrast, the AMF module avoids token-level pairwise interactions and instead performs interaction based on the overall confidence scores of the image and tabular modalities, rendering its complexity effectively independent of image dimensions. This advantage ensures the scalability and practicality of AMF in real-world applications involving high-resolution inputs.
4.5 Robustness to Incomplete Tabular Data
In the IG analysis of the supervised tabular models, age was identified as the most important feature in the OL3I dataset by FT-Mamba, FT-Transformer, and ResNet. Therefore, we masked the true age value, and a new $R_{conf}$ of 1.03 was established based on unimodal performance, as shown in Table A.1. The AUC of multimodal models using incomplete tabular data, along with their absolute decrease compared to the complete data condition, are shown in Table 4.
Table 4: AUCs of multimodal models on incomplete OL3I data and their absolute decrease compared to the complete setting.
[…]
AUC (%) Drop AUC (%) Drop
Concatenation - 70.02 $\pm$ 1.19 5.78 71.31 $\pm$ 2.20 5.18
Cross-attention - 72.58 $\pm$ 1.59 0.78 72.91 $\pm$ 0.50 0.57
AMF Module 0.95 74.88 $\pm$ 0.31 2.09 73.82 $\pm$ 0.84 4.54
AMF Module 1.03 75.07 $\pm$ 0.56 1.90 73.89 $\pm$ 0.78 4.47
In terms of AUC performance, the AMF module achieved the best results under incomplete tabular data conditions when $R_{conf}=1.03$, reaching 75.07% and 73.89% under frozen and trainable strategies, respectively—substantially outperforming other fusion methods. Regarding performance degradation, the cross-attention module exhibited the smallest AUC drop, suggesting a certain degree of robustness to incomplete tabular features. However, this small drop largely stems from its lower baseline performance when the tabular data is complete. Despite the smaller degradation, its absolute performance remains clearly inferior to that of the AMF module, indicating suboptimal fusion effectiveness. By contrast, the concatenation strategy suffered from over 5 percentage points of AUC drop in both missing conditions, reflecting a lack of effective compensation mechanisms in the presence of missing tabular inputs and thus poor robustness.
Further examining the response mechanism, the AMF module mitigates the negative impact of missing tabular features by explicit confidence control. Specifically, as the confidence in the tabular modality declines with increased feature missingness, the AMF module compensates by increasing $R_{conf}$ from 0.95 to 1.03, thereby encouraging the model to rely more heavily on the image modality. In contrast, the robustness of the cross-attention module primarily arises from implicit pattern fitting over large-scale data, lacking interpretability and generalizability in low-resource or high-missingness settings.
4.6 Ablation Studies on the AMF Module Design
4.6.1 Ablation Study on Modality Confidence Ratio
Table 5: AUCs of AMF-MedIT with different modality confidence ratios on the OL3I dataset.
$\mathbf{R_{conf}}$ $\mathbf{L_{img}}$ $\mathbf{L_{tab}}$ AUC (%)
[…]
None 2048 64 75.80 $\pm$ 0.24 76.49 $\pm$ 0.60
1.00 1024 1024 76.62 $\pm$ 0.90 78.33 $\pm$ 0.15
0.97 1009 1039 76.04 $\pm$ 1.09 78.08 $\pm$ 0.27
0.951 998 1059 76.97 $\pm$ 0.75 78.36 $\pm$ 0.12
0.10 342 1706 76.83 $\pm$ 0.37 78.16 $\pm$ 0.36
  • 1
    The value selected for FT-Mamba based on unimodal performance.
As shown in Table 5, the highest AUCs were achieved when $R_{conf}=0.95$, which aligns with the confidence ratio estimated from the unimodal performance of FT-Mamba and enables a balanced, task-consistent integration of both modalities. In contrast, settings based on mismatched modality pairs (e.g., $R_{conf}=0.97$ or $1.0$) resulted in performance degradation, suggesting that accurate allocation of unimodal feature dimensions is essential to maintain a balanced modality contribution and achieve effective fusion. In more extreme cases (e.g., $R_{conf}=0.1$ or None), the AUCs still declined compared to the optimal configuration, reinforcing that the adjustment of unimodal features is not arbitrary but grounded in meaningful modality-specific information.
These results collectively highlight that the effectiveness of the modulation-based fusion depends critically on appropriately scaling unimodal representations according to their task relevance. Therefore, it is necessary to introduce a modality confidence ratio based on actual modality reliability during fusion, which offers a simple yet interpretable way to achieve balanced unimodal feature scaling.
4.6.2 Ablation Study on Loss Functions
Table 6: AUCs of AMF-MedIT with different loss configurations on the OL3I dataset.
$\mathcal{L}_{\text{ce}}$ $\mathcal{L}_{\text{leakage}}$ $\mathcal{L}_{\text{density}}$ AUC (%)
[…]
$\times$ $\times$ 73.11$\pm$3.99 78.07$\pm$0.64
$\times$ 75.46$\pm$1.71 78.05$\pm$0.36
$\times$ 74.11$\pm$3.49 78.08$\pm$0.35
76.97$\pm$0.75 78.36$\pm$0.12
The AUCs under different loss configurations are presented in Table 6. The best performance is achieved when all three losses are jointly applied. Removing the leakage loss $\mathcal{L}_{\text{leakage}}$ leads to a substantial drop in performance, especially under the frozen strategy (-2.86% AUC), indicating that the absence of information flow regulation across modalities causes significant semantic leakage and degraded fusion quality. Similarly, omitting the density loss also reduces performance and introduces instability, as it fails to constrain the feature magnitude alignment between the image and tabular modalities, which can impede effective fusion.
The improvement brought by these two auxiliary losses is more evident in the frozen training setting, where the fusion module receives no gradient from the backbone. In such cases, relying solely on the classification loss $\mathcal{L}_{\text{ce}}$ is insufficient for guiding the training of the AMF module, leading to higher performance variance and lower overall accuracy. These results collectively validate the necessity of both density and leakage losses for effective multimodal training, as they enhance representation compatibility, stabilize optimization, and ultimately improve classification performance.
5 Discussion
In this work, we propose AMF-MedIT, an align-modulation-fusion framework for medical image and tabular data. Our motivation stems from the significant cross-modal differences in dimension and confidence levels during fusion, which are further exacerbated by the scarcity of medical data. To address this, we propose the AMF module, a simple yet efficient solution for multimodal fusion. The AMF module not only achieves effective multimodal classification performance on both medical and natural datasets, but also offers high computational efficiency and strong adaptability to incomplete tabular data. Furthermore, for the noisy medical tabular data, we present FT-Mamba for tabular feature extraction. It achieves a favorable balance between data efficiency and feature discriminability and plays a pivotal supervisory role in image-tabular contrastive pretraining.
5.1 Discussion of Fusion Methods
In the study of fusion methods, we compared our proposed AMF module with two widely adopted approaches—naive feature concatenation and the cross-attention mechanism—across three key aspects: performance, computational efficiency, and robustness to missing tabular data. Naive concatenation is a straightforward and commonly used approach in medical multimodal learning. However, its fixed and non-interactive network structure lacks the capacity to regulate the dimensionality and contribution of each modality. As a result, compact tabular features are placed at a disadvantage, leading to suboptimal fusion performance. Furthermore, the rigidity of this approach makes it less robust to variations in downstream data distributions, limiting its adaptability in real-world clinical scenarios.
The cross-attention module, on the other hand, takes multimodal token sequences with equal dimensions as input. By leveraging explicit token-level attention, it enables deeper cross-modal interaction and demonstrates strong robustness in handling incomplete tabular data. Nevertheless, its full attention architecture introduces high computational complexity and memory overhead, which may hinder its applicability in data-limited medical contexts or real-time clinical deployments.
To strike a better balance between effectiveness and efficiency, the AMF module replaces token-level attention with a lightweight, confidence-driven modulation strategy. Rather than enforcing equal or fixed feature dimensions, AMF adaptively adjusts the unimodal feature dimensions according to its global confidence. In addition, we propose a density loss that serves as a soft normalization mechanism to align the density of features across modalities. The adjustments in both dimension and density help mitigate scale mismatches between heterogeneous modalities and facilitate the formation of a unified cross-modal representation space. Moreover, our design incorporates a feature mask along with a leakage loss, which permits flexible information flow and compression, enabling the unified space to preserve complementary modality-specific information. This unified yet relaxed-constrained representation space reconciles the representational and confidence discrepancies between tabular and image modalities, supporting effective multimodal integration as confirmed by our experimental results. At the same time, the module’s simplicity allows for efficient fine-tuning with limited data and fast inference, making it practical for real-world deployment.
As another simplification, we introduce a scalar modality confidence ratio $R_{conf}$ which is derived from preliminary empirical results. The incorporation of prior knowledge is particularly valuable in the medical domain. On the one hand, it helps mitigate data scarcity by leveraging external expertise and reduces the reliance on purely data-driven learning. On the other hand, it enables the integration of medical knowledge or clinical experience into future automated systems, enhancing the interpretability of the fusion process and its acceptance among medical professionals. Beyond facilitating effective fusion, $R_{conf}$ also enhances the model’s adaptivity to various clinical scenarios, including tabular data missingness in our experiments, image quality degradation, and even modality absence. Importantly, AMF achieves this with fewer parameters and orders-of-magnitude lower computational cost than cross-attention, making it more suitable for data-scarce, latency-sensitive, or resource-constrained medical applications.
5.2 Discussion of Tabular Encoders for Medical Data
In the study of tabular encoders, we compared the performance of four representative architectures—MLP, ResNet, FT-Transformer, and FT-Mamba—and further conducted interpretability analyses to understand their working mechanisms and roles in contrastive learning. As a kind of structural text, medical tabular data also exhibits sequential characteristics. For such data, sequence models like FT-Mamba and FT-Transformer demonstrate superior feature discrimination through their respective selection and self-attention mechanisms. This ability to automatically identify relevant features from a large pool of redundant variables is particularly important in processing EHRs, which encompass a wide range of diseases, time spans, and heterogeneous variables. Enhanced feature discrimination not only reduces the need for manual feature engineering but also enables the model to focus on clinically relevant patterns, potentially uncovering latent medical associations. However, among these two sequence models, FT-Transformer’s reliance on full attention matrices leads to overfitting on small datasets, while FT-Mamba is more efficient in low-data settings by selectively focusing on relevant information.
Due to its advantages, FT-Mamba provides effective supervision for imaging modality during contrastive pretraining, guiding the encoder to flexibly focus on multiple key regions. In contrast, ResNet exhibits strong locality in tabular feature extraction, showing a marked focus on a single feature—age—which induces a singular and coarse focus in the image modality. The visual attention is primarily concentrated on the lumbar spine, likely because the correlation between patient age and bone density (Jiang et al., 2022) is implicitly leveraged by the model during contrastive learning. Our interpretability analysis reveals that tabular encoders with different representational capacities exert distinctly different supervisory effects on the image modality during contrastive learning. Therefore, the selection of the tabular encoder plays a crucial role in multimodal pretraining, as it directly influences the representational power of the learned features.
5.3 Limitations and Future Directions
Our study has several limitations. First, the results on the DVM dataset indicate that FT-Mamba still underperforms FT-Transformer on large-scale data. While FT-Mamba serves as a compromise for limited medical data, we expect FT-Transformer to perform better as large-scale medical datasets become more common. Second, the $R_{conf}$ configuration in the AMF module is sensitive, with slight variations significantly affecting performance in our experiments. Third, our experiments were limited to two multimodal datasets, and further validation on a broader range of medical datasets is needed. Finally, our model is restricted to classification tasks, and other medical image-tabular applications, such as segmentation and report generation, require further exploration.
Future research could investigate automated optimization of $R_{conf}$ based on dataset characteristics and develop guidelines informed by domain experts to improve its robustness. Additionally, evaluating our approach on diverse datasets, including those from different imaging modalities like magnetic resonance imaging and ultrasound, would enhance its generalizability. Extending the AMF module to support segmentation tasks could also broaden its applicability in medical research.
6 Conclusion
In conclusion, we proposed AMF-MedIT, an efficient Align-Modulation-Fusion framework for integrating medical image and tabular data, to address the cross-modal disparities during fusion and noisy medical tabular inputs. At its core, the AMF module leverages streamlined modulation and prior knowledge to effectively coordinate cross-modal disparities, providing an accurate and adaptable fusion method for various downstream conditions. Additionally, FT-Mamba enhances the performance of the tabular encoder, achieving an optimal balance between performance and data efficiency. Furthermore, our interpretability study on image-tabular contrastive learning reveals the impact of encoder selection on multimodal pretraining. Our approach enhances data efficiency and mitigates reliance on large annotated datasets, offering a novel solution to challenges in medical multimodal fusion. We hope these findings will inspire advancements in medical multimodal models and their applications in healthcare.
CRediT Authorship Contribution Statement
Congjing Yu: Conceptualization, Data curation, Methodology, Validation, Writing - original draft, Writing - review & editing. Jing Ye: Conceptualization, Validation, Writing - review & editing. Yang Liu: Conceptualization, Validation, Writing - review & editing. Xiaodong Zhang: Conceptualization, Supervision. Zhiyong Zhang: Conceptualization, Funding acquisition, Supervision, Writing - review & editing.
or transparency, we require corresponding authors to provide co-author contributions to the manuscript using the relevant CRediT roles. The CRediT taxonomy includes 14 different roles describing each contributor’s specific contribution to the scholarly output. The roles are: Conceptualization; Data curation; Formal analysis; Funding acquisition; Investigation; Methodology; Project administration; Resources; Software; Supervision; Validation; Visualization; Roles/Writing - original draft; and Writing - review & editing. Note that not all roles may apply to every manuscript, and authors may have contributed through multiple roles.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgments
The work was supported by the Science and Technology Planning Project of Key Laboratory of Advanced IntelliSense Technology, Guangdong Science and Technology Department (2023B1212060024).
Appendix A Implementation Details
A.1 Data Preprocessing
For CT images in the OL3I dataset, the extreme grayscale values (greater than 275 and less than -125) were removed to enhance the image contrast. Then, the pixel values of images from both datasets were normalized to the range of 0 to 1.
For missing tabular data in the OL3I dataset, we applied median imputation. For tabular data in the DVM dataset, 13 features including 9 continuous features (Adv_year, Adv_month, Reg_year, Runned_Miles, Price, Seat_num, Door_num, Entry_price, Engine_size) and 4 categorical features (Color, Bodytype, Gearbox, Fuel_type) were remained for input. The continuous tabular features in both datasets were normalized using z-score normalization before being fed into the model.
A.2 Multimodal Pretraining
The pretraining procedures and augmentations followed those described in (Hager et al., 2023), where image augmentations were applied with a probability of 0.95 and tabular features were corrupted at a rate of 0.3. The optimizer used was Adam with a weight decay of $1e^{-4}$. The learning rate was selected based on validation performance, with possible values of $(3e^{-3},3e^{-4},3e^{-4})$. The model was trained for 500 epochs using a cosine annealing scheduler with a warmup period of 10 epochs.
A.3 Supervised Training and Fune-tuning
Considering the severe class imbalance in the OL3I dataset, where class 0 contains 7,784 samples and class 1 only 355 samples (a ratio of approximately 20:1), the weight factor $\alpha$ in the weighted CE loss was set to 0.1 for class 0 and 2 for class 1. Additionally, a weighted sampler strategy was applied to the OL3I dataset, where class 0 in the training set was downsampled by a factor of 0.5, and class 1 was oversampled by a factor of 5.
To determine the embedding dimension $D_{tab}$ for tabular encoders, we evaluated multiple candidate dimensions during supervised training on the OL3I dataset. For MLP and ResNet, the values tested were $\{128,256,512,1024\}$, and for FT-Transformer and FT-Mamba, $\{32,64,96,128,256,384\}$ were considered. During the fine-tuning of the AMF module, $\lambda_{1}$ and $\lambda_{2}$ were empirically set to 5.0 to ensure that the density and leakage losses have a comparable magnitude and regularization strength to the cross-entropy loss. To optimize the training hyperparameters, we tested three learning rates $\{1e^{-2},1e^{-3},1e^{-4}\}$ using a fixed seed of 2022 for all supervised and fine-tuning experiments. In the fine-tuning of the cross-attention module, we observed severe overfitting and thus tested weight decay values of $\{1e^{-1},1e^{-2},1e^{-3},1e^{-4}\}$ to regularize training. In all other cases, the weight decay factor was fixed to 0. The optimal values were selected based on the validation performance.
All experiments utilized the Adam optimizer. Early stopping was applied based on the validation metric, triggered if no improvement greater than 0.0002 was observed over 10 consecutive epochs. After determining the optimal hyperparameters, each model was retrained using five random seeds (2022–2026) to mitigate randomness.
A.4 Quantifying the Modality Confidence Ratio
In our experiments, we assessed the confidence levels of unimodal feature vectors, $\textbf{Confidence}(\boldsymbol{z}_{img})$ and $\textbf{Confidence}(\boldsymbol{z}_{tab})$, based on their respective unimodal performance. Specifically, after multimodal pretraining, we constructed unimodal models by utilizing the pretrained encoders and fine-tuned them on downstream tasks with a frozen strategy. The modality confidence ratio $R_{\text{conf}}$ was then calculated as the ratio of the image-to-tabular unimodal model metrics on the testing set. The unimodal performance and the corresponding modality confidence ratio $R_{\text{conf}}$ are presented in Table A.1.
Table A.1: Unimodal performance of pretrained encoders and corresponding modality confidence ratio.
Image Encoder Tabular Encoder OL3I AUC (%) OL3I DVM ACC (%) DVM
Image Tabular ${R_{conf}}$ Image Tabular ${R_{conf}}$
ResNet-50 MLP 74.99 77.41 0.97 66.11 79.65 0.83
ResNet 74.99 75.64 1.00 45.23 77.95 0.58
FT-Transformer 75.62 74.62 1.01 34.20 71.23 0.48
FT-Mamba 72.92 76.21 0.95 60.20 66.78 0.90
FT-Mamba1 72.92 70.98 1.03 - - -
  • 1
    Perfomance on incomplete tabular data.
Appendix B Detailed Results of Integrated Gradient Analysis
Table B.1 presents the top 10 most important tabular features and their corresponding importance scores as identified by the IG analysis for the four unimodal tabular models. While MLP, ResNet, and FT-Transformer tend to focus on several variables that are less relevant to IHD, FT-Mamba demonstrates a strong emphasis on features closely associated with the cardiovascular system.
Table B.1: Top 10 most important tabular features for MLP, ResNet, FT-Transformer and FT-Mamba unimodal tabular models.
Tabular Model Feature Description Importance score
MLP gender Whether the patient is male 0.056
age_at_scan Age at time of abdominal scan in years 0.027
Evaluation_14 Number of times a CPT code was documented in the group: Preventive Medicine Services 0.022
V04 Number of prescriptions that match ATC code classification: Diagnostic Agents 0.021
smoker Whether it was documented that the patient is a smoker in the year prior to abdominal scan 0.021
exp_wt_lab_diastolic Exponentially weighted average of measurements diastolic blood pressure during the year prior to abdominal scan [mm Hg] 0.018
S02 Number of prescriptions that match ATC code classification: Otologicals 0.018
D10 Number of prescriptions that match ATC code classification: Anti-Acne Preparations 0.017
B01 Number of prescriptions that match ATC code classification: Antithrombotic Agents 0.017
A04 Number of prescriptions that match ATC code classification: Antiemetics And Antinauseants 0.016
ResNet age_at_scan Age at time of abdominal scan in years 0.180
CCI Charlson Comorbidity Index 0.034
B01 Number of prescriptions that match ATC code classification: Antithrombotic Agents 0.021
J02 Number of prescriptions that match ATC code classification: Antimycotics For Systemic Use 0.016
Chapter_II_4 Number of times an ICD was documented in the year prior to imaging: Neoplasms of uncertain or unknown behaviour 0.014
C09 Number of prescriptions that match ATC code classification: Agents Acting On The Renin-Angiotensin System 0.014
Medical_9 Number of times a CPT code was documented in the group: Special Otorhinolaryngologic Services and Procedures 0.013
Evaluation_14 Number of times a Current Procedural Terminology code was documented in the group: Preventive Medicine Services 0.013
M02 Number of prescriptions that match ATC code classification: Topical Products For Joint And Muscular Pain 0.013
exp_wt_lab_systolic Exponentially weighted average of measurements of systolic blood pressure during the year prior to abdominal scan [mm Hg] 0.012
FT-Transformer age_at_scan Age at time of abdominal scan in years 0.362
CCI Charlson Comorbidity Index 0.322
C08 Number of prescriptions that match ATC code classification: Calcium Channel Blockers 0.098
A12 Number of prescriptions that match ATC code classification: Mineral Supplements 0.057
J05 Number of prescriptions that match ATC code classification: Antivirals For Systemic Use 0.027
gender Whether the patient is male 0.026
A06 Number of prescriptions that match ATC code classification: Drugs For Constipation 0.011
Surgery_1 Number of times a CPT code was documented in the group: General Surgical Procedures 0.010
C10 Number of prescriptions that match ATC code classification: Lipid Modifying Agents 0.010
Evaluation_14 Number of times a CPT code was documented in the group: Preventive Medicine Services 0.010
FT-Mamba age_at_scan Age at time of abdominal scan in years 0.452
CCI Charlson Comorbidity Index 0.085
gender Whether the patient is male 0.063
C03 Number of prescriptions that match ATC code classification: Diuretics 0.032
Evaluation_14 Number of times a CPT code was documented in the group: Preventive Medicine Services 0.013
C09 Number of prescriptions that match ATC code classification: Agents Acting On The Renin-Angiotensin System 0.012
B05 Number of prescriptions that match ATC code classification: Blood Substitutes And Perfusion Solutions 0.011
exp_wt_lab_systolic Exponentially weighted average of measurements of systolic blood pressure during the year prior to abdominal scan [mm Hg] 0.010
V06 Number of prescriptions that match ATC code classification: General Nutrients 0.009
Chapter_IX_3 Number of times an ICD was documented in the year prior to imaging: Hypertensive diseases 0.008